# Strong and weak wave turbulence regimes in Bose-Einstein condensates

### v2

@@ Abstract
When a turbulent Bose-Einstein condensate is driven out-of-equilibrium at a scale much smaller than the system size, nonlinear wave interactions transfer particles towards large scales in an inverse cascade process. In this work, we study numerically wave turbulence in a three-dimensional Bose-Einstein condensate in forced and dissipated inverse cascade settings. We observe that when the forcing rate increases, thereby increasing the particle flux, the turbulence spectrum gradually transitions from the weak-wave Kolmogorov-Zakharov cascade to a critical balance state characterized by a range of scales with balanced linear and nonlinear dynamic timescales. Further forcing increases lead to a coherent condensate component superimposed with Bogoliubov-type acoustic turbulence. We then use our predictions and numerical data to formulate a new out-of-equilibrium equation of state for the $3D$ inverse cascade.
@@
respectively. In a turbulent setting, energy is transferred towards small scales (direct cascade), whereas particles exhibit an inverse cascade towards large scales. Most of turbulent BEC experiments have studied the direct energy cascade Navon et al. (2019); Gałka et al. (2022), whereas inverse particle turbulent transfers have been recently observed both in 2D and 3D experiments Karailiev et al. (2024); Moreno-Armijos et al. (2025).
In this Letter, we focus on the inverse particle cascade in 3D. In particular, we study how the system transitions from weak to strong wave turbulence when the particle flux $|Q_{0}|$ increases. The advantage of this setting is that the WWT provides an exact and simple analytical prediction, unlike the 3D direct energy cascade or the cascades in 2D BECs that present some mathematical issues.
As usual in turbulence, let us start by dimensional considerations. The main observable in the turbulent cascade is the waveaction spectrum, $n_{\bf k}(t)\equiv n({\bf k},t)=\frac{V}{(2\pi)^{3}}\langle|\hat{\psi}_{\bf k}(t)|^{2}\rangle,$ where $\hat{\psi}_{\bf k}=\frac{1}{L^{3}}\int\psi(\mathbf{x})e^{i{\bf{k}}\cdot\mathbf{x},}\,\mathrm{d}\,\mathbf{x}$ is the Fourier transform of $\psi$, and the brackets denote the average over fluctuations. Note that $n_{\bf k}$ is dimensionless, and from Eq. (1), $[\alpha]=length^{2}/time$ and $[\beta]=length^{3}/time$. As $[Q_{0}]=length^{-3}/time$, then by matching dimensions we have
@@
where $x$ and $y$ are a priori arbitrary exponents, and the constant $C$ and the two terms in parentheses are dimensionless. Here, $x$ fixes the dependence of the waveaction spectrum on scales whereas $y$ determines the behavior of the so-called out-of-equilibrium equation of state Dogra et al. (2023); Zhu et al. (2024). Indeed, the the spectrum’s amplitude dependence on the flux $Q_{0}$ is analogous to the particle number dependence on the chemical potential at thermal equilibrium. The WWT predicts the Kolmogorov-Zakharov (KZ) spectrum that fixes all the exponents, including the universal dimensionless constant $C$ Zhu et al. (2023):
@@
One of the main hypotheses of the WWT, is the assumption of random phases and amplitudes. When the nonlinearity is strong, phases might be strongly correlated and spectrum could be guessed by invoking the critical balance (CB) assumption where linear and nonlinear-transfer terms in (1) are balanced. It follows that the spectrum is independent of the flux ($y=0$). Such an assumption, first evoked in the context of water waves, led to the so-called Phillips spectrum. Here, unlike water waves where the CB assumption fixes all exponents, we still need to fix the scaling exponent $x$. For GPE, the CB spectrum was first introduced in Proment et al. (2009) using the GPE (1) written in Fourier space
@@
where $\sum_{123}$ stands for a sum over all wave vectors ${\bf{k}}_{1},{\bf{k}}_{2},{\bf{k}}_{3}$, $\psi_{\bf i}=\psi({\bf{k}}_{i})$, and $\delta^{{\bm{k}}1}_{23}$ is a Kronecker-$\delta$: $1$ if ${\bf{k}}+{\bf{k}}_{1}={\bf{k}}_{2}+{\bf{k}}_{3}$, 00 otherwise. Note that for both WWT and CB, the leading-order nonlinear effect in (6) is provided by the diagonal terms in the sum, i.e., the ones with ${\bf{k}}={\bf{k}}_{2},{\bf{k}}_{1}={\bf{k}}_{3}$ and ${\bf{k}}={\bf{k}}_{3},{\bf{k}}_{1}={\bf{k}}_{2}$. They lead to a nonlinear frequency shift $\omega_{k}\to\omega_{k}+2N$, but do not affect the evolution of $|\hat{\psi}_{{\bm{k}}}|$ and, therefore, do not lead to any spectral redistribution. They can be eliminated by passing into the rotating frame, $\hat{\psi}_{\bf k}\to e^{-2iNt}\hat{\psi}_{\bf k}$. Applying the CB argument to (6), we find that $\omega_{k}|\hat{\psi}_{\bf k}|\sim\beta|\hat{\psi}_{\bf k}|^{3}(k/k_{\rm cor})^{6}$ where $k_{\rm cor}$ is the correlation distance in the $k$-space (it is vanishingly small in the WWT). The $(k/k_{\rm cor})^{6}$-term estimates the number of modes interacting with mode $k$. With $n_{\bf k}\sim|\hat{\psi}_{\bf k}|^{2}/k_{\rm cor}^{3}$, this leads to $n_{\bf k}\sim\frac{\alpha}{\beta k}\frac{k_{\rm cor}^{3}}{k^{3}}.$ In the limit $L\to\infty$, this expression should be independent of $L$, which arises if we postulate $k_{\rm cor}\sim\alpha/\beta$. This finally gives
@@
Numerical simulations.— We perform numerical simulations of the forced-dissipated GPE using the standard massively-parallel pseudo-spectral code FROST Krstulovic (2020) with a fourth-order Exponential Runge-Kutta temporal scheme (see Zhu et al. (2023)). We set $L=2\pi$, $\alpha=\beta=1$, and use grids of $N_{p}^{3}$ collocation points, with $N_{p}=512$. We add a forcing term $F_{\bf k}(t)$ and a dissipation term $-D_{\bf k}\widehat{\psi}_{\bf{k}}(t)$ to the RHS of the Fourier-space GPE (6). To obtain inverse cascade the forcing term is supported on a narrow band at high $k$’s around the forcing wavenumber $k_{\rm f}$ for $k\in[k_{\rm f}-1,k_{\rm f}+1]$, and it is given by $\mathrm{d}F_{\bf k}(t)=f_{0}\mathrm{d}{W}_{\bf k}$, where $W_{\bf k}$ is the Wiener process, and the parameter $f_{0}$ controls the amplitude of the forcing. Dissipation is of the form $D_{\bf k}=D_{\rm L}k^{-r}$ at the largest scales ($0<k\leq k_{\rm L}\ll k_{\rm f}$), and $D_{\bf k}=D_{\rm R}k^{s}$ at the very small scales less than the forcing scales ($k_{\rm f}+1<k\leq k_{\rm max}$). Additionally, the condensate mode ($k=0$) is dissipated at a constant rate $D_{\bf 0}$. We optimize the parameters of forcing and dissipation to enlarge the inertial range for a fixed resolution, ensuring that the simulations that are well-resolved and minimizing bottlenecks at the dissipation scales. All the numerical parameters are listed in Appendix A. We perform simulations with different values of the forcing strength $f_{0}$ in order to study the transition from the weak to the strong WT regimes. The $k$-space particle flux $Q(k)$ is computed directly from the RHS of the GPE (6) as in Zhu et al. (2023). We evaluate $Q_{0}$ by measuring the value of $Q(k)$ in the inertial range where it is approximately constant. Note that $f_{0}$ and $|Q_{0}|$ are monotonically related, so we discuss our results in terms of the strength of the particle flux. Finally, our numerical measurements are performed when a statistically steady state is reached.
@@
We will now discuss the out-of-equilibirum equation of state Dogra et al. (2023); Zhu et al. (2024), which relates the global amplitude of the waveaction spectrum to the input flux, in our case $|Q_{0}|$. In practice, we fit our data with $n_{k}=Ak^{-x}$ in the different observed ranges. The exponents $x$ was found to be very close those discussed earlier, and the amplitude $A$ is strongly dependent on $Q_{0}$. Readers can find the fitting method and fitting exponents $x$ in the Appendix C. Figure 3 shows the plot of the spectrum’s amplitude $A$ vs. the particle flux in the inertial range $|Q_{0}|$ for our runs covering the whole range of particle fluxes, from low to high (the flux functions $Q(k)$ for all the runs are shown in Appendix B). For low-flux runs, we observe $A\sim|Q_{0}|^{1/3}$ in agreement with the weak four-wave WT. For intermediate fluxes, the spectrum at the high-$k$ part remains weak and still follows $A\sim|Q_{0}|^{1/3}$. The low-$k$ part, which exhibits the CB exponent (see Appendix C), displays some dependence on the flux that could be fitted by $A\sim|Q_{0}|^{2/3}$. Such an exponent evokes some type of hydrodynamic Kolmogorov turbulence, but we lack a theoretical explanation and scaling range is too narrow. Finally, for high-flux runs, we only measure $A$ for the large-$k$ thermodynamic part because the low-$k$ condensate spectrum range is too short for a reliable fitting. Here we see $A=T\sim|Q_{0}|^{1/2}$, which fulfills our prediction for the temperature scaling of the three-wave process.
Figure 3: “Equation of state” plot covering the weak, intermediate, and strong flux regimes.
@@ Acknowledgements.
This work was funded by the Simons Foundation Collaboration grant Wave Turbulence (Award ID 651471). This work was granted access to the high-performance computing facilities under GENCI (Grand Equipement National de Calcul Intensif) A0102A12494 (IDRIS and ADASTRA), the OPAL infrastructure from Université Côte d’Azur, supported by the French government, through the UCAJEDI Investments in the Future project managed by the National Research Agency (ANR) under Reference No. ANR-15-IDEX-01.
@@ Appendix C Method to generate the EoS figure
Suppose the spectrum is of the form $n=A\,k^{-x}$, we compute the values of $A$ and $x$ by fitting the power-law range of the numerically observed spectrum with $\ln{n}=\ln{A}-x\ln{k}$. Figure S2
presents how we fit spectra for the strong turbulence cases. Refer to caption
@@
The fitting exponents corresponding to Figure 3 is shown in Figure S3.
Figure S3: The fitting exponents correspond to Figure 3. The theoretical predictions for weak wave turbulence, critical balance, and the Bogoliubov spectrum are $-7/3$, $-4$, and $-2$, respectively.

### v4

@@ Abstract
When a turbulent Bose-Einstein condensate is driven out-of-equilibrium at a scale much smaller than the system size, nonlinear wave interactions transfer particles towards large scales in an inverse cascade process. In this work, we study numerically wave turbulence in a three-dimensional Bose-Einstein condensate in forced and dissipated inverse cascade settings. We observe that when the forcing rate increases, thereby increasing the particle flux, the turbulence spectrum gradually transitions from the weak-wave Kolmogorov-Zakharov cascade to a critical balance state characterized by a range of scales with balanced linear and nonlinear dynamic timescales. Further forcing increases lead to a coherent condensate component superimposed with Bogoliubov-type acoustic turbulence. The role of vortices in such a strongly forced state is marginal, which makes this new state very different from the strongly turbulent state composed from a tangle of quantized vortex lines. We then use our predictions and numerical data to formulate a new out-of-equilibrium equation of state for the $3D$ inverse cascade.
@@
respectively. In a turbulent setting, energy is transferred towards small scales (direct cascade), whereas particles exhibit an inverse cascade towards large scales. Most of turbulent BEC experiments have studied the direct energy cascade Navon et al. (2019); Gałka et al. (2022), whereas inverse particle turbulent transfers have been recently observed both in 2D and 3D experiments Karailiev et al. (2024); Moreno-Armijos et al. (2025). (The inverse particle cascade should not be confused with the inverse energy cascades and clustering in the 2D BEC quantized vortex turbulence studied in Johnstone et al. (2019); Gauthier et al. (2019).)
In this Letter, we focus on the inverse particle cascade in 3D. In particular, we study how the system transitions from weak to strong wave turbulence when the particle flux per unit of volume $|Q_{0}|$ increases. The advantage of this setting is that the WWT provides an exact and simple analytical prediction, unlike the 3D direct energy cascade or the cascades in 2D BECs that present some mathematical issues.
As usual in turbulence, let us start by dimensional considerations. The main observable in the turbulent cascade is the waveaction spectrum, $n_{\bf k}(t)\equiv n({\bf k},t)=\frac{V}{(2\pi)^{3}}\langle|\hat{\psi}_{\bf k}(t)|^{2}\rangle,$ where $\hat{\psi}_{\bf k}=\frac{1}{L^{3}}\int\psi(\mathbf{x})e^{i{\bf{k}}\cdot\mathbf{x},}\,\mathrm{d}\,\mathbf{x}$ is the Fourier transform of $\psi$, and the brackets denote the average over fluctuations. Note that $n_{\bf k}$ is dimensionless, and from Eq. (1), $\alpha]=length^{2}/time$ and $\beta]=length^{3}/time$. As $Q_{0}]=length^{-3}/time$, then by matching dimensions we have
@@
where $C$ is a dimensionless constant, the two terms in parentheses are independent dimensionless combinations and we assumed a power law dependence on these parameters with $x$ and $y$ being a priori arbitrary exponents. Here, $x$ fixes the dependence of the waveaction spectrum on scales whereas $y$ determines the behavior of the so-called out-of-equilibrium equation of state Dogra et al. (2023); Zhu et al. (2024). Indeed, the the spectrum’s amplitude dependence on the flux $Q_{0}$ is analogous to the particle number dependence on the chemical potential at thermal equilibrium. The WWT predicts the Kolmogorov-Zakharov (KZ) spectrum that fixes all the exponents, and the universal dimensionless constant $C$ Zhu et al. (2023):
@@
One of the main hypotheses of the WWT, is the assumption of random phases and amplitudes. When the nonlinearity is strong, phases might be strongly correlated and spectrum could be guessed by invoking the critical balance (CB) assumption where linear and nonlinear-transfer terms in (1) are balanced. It follows that the spectrum is independent of the flux ($y=0$). Such an assumption, first evoked in the context of water waves, led to the so-called Phillips spectrum Phillips (1958). Here, unlike water waves where the CB assumption fixes all exponents, we still need to fix the scaling exponent $x$. For GPE, the CB spectrum was first introduced in Proment et al. (2009) using the GPE (1) written in Fourier space
@@
where $\sum_{123}$ stands for a sum over all wave vectors ${\bf{k}}_{1},{\bf{k}}_{2},{\bf{k}}_{3}$, $\psi_{\bf i}=\psi({\bf{k}}_{i})$, and $\delta^{{\bm{k}}1}_{23}$ is a Kronecker-$\delta$: 11 if ${\bf{k}}+{\bf{k}}_{1}={\bf{k}}_{2}+{\bf{k}}_{3}$, 0 otherwise. Note that for both WWT and CB, the leading-order nonlinear effect in (6) is provided by the diagonal terms in the sum, i.e., the ones with ${\bf{k}}={\bf{k}}_{2},{\bf{k}}_{1}={\bf{k}}_{3}$ and ${\bf{k}}={\bf{k}}_{3},{\bf{k}}_{1}={\bf{k}}_{2}$. They lead to a nonlinear frequency shift $\omega_{k}\to\omega_{k}+2\beta N$, but do not affect the evolution of $|\hat{\psi}_{{\bm{k}}}|$ and, therefore, do not lead to any spectral redistribution. They can be eliminated by passing into the rotating frame, $\hat{\psi}_{\bf k}\to e^{-2i\beta Nt}\hat{\psi}_{\bf k}$. Applying the CB argument to (6), we find that $\omega_{k}|\hat{\psi}_{\bf k}|\sim\beta|\hat{\psi}_{\bf k}|^{3}(k/k_{\rm cor})^{6}$ where $k_{\rm cor}$ is the correlation distance in the $k$-space (it is vanishingly small in the WWT). The $(k/k_{\rm cor})^{6}$-term estimates the number of modes interacting with mode $k$. With $n_{\bf k}\sim|\hat{\psi}_{\bf k}|^{2}/k_{\rm cor}^{3}$, this leads to $n_{\bf k}\sim\frac{\alpha}{\beta k}\frac{k_{\rm cor}^{3}}{k^{3}}.$ In the limit $L\to\infty$, this expression should be independent of $L$, which arises if we postulate $k_{\rm cor}\sim\alpha/\beta$. This finally gives
@@
Numerical simulations.— We perform numerical simulations of the forced-dissipated GPE using the standard massively-parallel pseudo-spectral code FROST Krstulovic (2020) with a fourth-order Exponential Runge-Kutta temporal scheme (see Zhu et al. (2023)). We set $L=2\pi$, $\alpha=\beta=1$, and use grids of $N_{p}^{3}$ collocation points, with $N_{p}=512$. We add a forcing term $F_{\bf k}(t)$ and a dissipation term $D_{\bf k}\widehat{\psi}_{\bf{k}}(t)$ to the RHS of the Fourier-space GPE (6). To obtain inverse cascade the forcing term is supported on a narrow band at high $k$’s around the forcing wavenumber $k_{\rm f}$ for $k\in[k_{\rm f}-1,k_{\rm f}+1]$, and it is given by $\mathrm{d}F_{\bf k}(t)=f_{0}\mathrm{d}{W}_{\bf k}$, where $W_{\bf k}$ is the Wiener process, and the parameter $f_{0}$ controls the amplitude of the forcing. Dissipation is of the form $D_{\bf k}=D_{\rm L}k^{-r}$ at the largest scales ($0<k\leq k_{\rm L}\ll k_{\rm f}$), and $D_{\bf k}=D_{\rm R}k^{s}$ at the very small scales less than the forcing scales ($k_{\rm f}+1<k\leq k_{\rm max}$). Additionally, the condensate mode ($k=0$) is dissipated at a constant rate $D_{\bf 0}$. We optimize the parameters of forcing and dissipation to enlarge the inertial range for a fixed resolution, ensuring that the simulations that are well-resolved and minimizing bottlenecks at the dissipation scales. All the numerical parameters are listed in the Supplemental Material (SM). We perform simulations with different values of the forcing strength $f_{0}$ in order to study the transition from the weak to the strong WT regimes. The $k$-space particle flux $Q(k)$ is computed directly from the RHS of the GPE (6) as in Zhu et al. (2023). We evaluate $Q_{0}$ (the rate of particle loss per unit volume and unit time through low-k dissipation) by measuring the value of $Q(k)$ in the inertial range where it is approximately constant. Note that $f_{0}$ and $|Q_{0}|$ are monotonically related, so we discuss our results in terms of the strength of the particle flux. Finally, our numerical measurements are performed when a statistically steady state is reached.
@@
Let us clarify about the role of vortices in our system for different forcing intensities. Vortices are associated with incompressible kinetic energy, which arises from the Helmholtz decomposition of the regularized velocity field into solenoidal and potential parts. Fig. 3(a) displays the energy components for various values of $|Q_{0}|$: incompressible and compressible kinetic energies, internal energy, and quantum energy. For low $|Q_{0}|$, quantum energy is dominant and internal energy is minimal, consistent with weak nonlinearity. Compressible and incompressible energies are equal, and each is half the quantum energy. The presence of incompressible kinetic energy indicates vortices, but these are ”ghost” vortices without dynamical significance—they appear even in purely linear fields Berry and Dennis (2000). Interestingly, increasing $|Q_{0}|$ reduces incompressible kinetic energy to about 6–8% in the strongest forcing cases, reflecting a sparse distribution of small vortex loops (with radii of the order of the healing length). This state contrasts with strong superfluid hydrodynamic turbulence states consisting of polarized tangles of fully formed quantized vortex lines of all sizes Polanco et al. ; Fischer and Bradley (2025). In our system, strong vortex lines shrink and annihilate due to effective friction from the intense acoustic component. (Persistent vortex tangles would require sound damping.) Thus, the strong turbulence state reported here is dominated by interacting acoustic waves rather than dynamically significant hydrodynamic vortices.
We will now discuss the out-of-equilibirum equation of state Dogra et al. (2023); Zhu et al. (2024), which relates the global amplitude of the waveaction spectrum to the input flux, in our case $|Q_{0}|$. In practice, we fit our data with $n_{k}=Ak^{-x}$ in the different observed ranges. The exponents $x$ was found to be very close those discussed earlier, and the amplitude $A$ is strongly dependent on $Q_{0}$. Readers can find the fitting method and fitting exponents $x$ in the SM. Figure 3(b) shows the plot of the spectrum’s amplitude $A$ vs. the particle flux in the inertial range $|Q_{0}|$ for our runs covering the whole range of particle fluxes, from low to high (the flux functions $Q(k)$ for all the runs are shown in the SM). For low-flux runs, we observe $A\sim|Q_{0}|^{1/3}$ in agreement with the weak four-wave WT. For intermediate fluxes, the spectrum at the high-$k$ part remains weak and still follows $A\sim|Q_{0}|^{1/3}$. The low-$k$ part, which exhibits the CB exponent (see SM), displays some dependence on the flux that could be fitted by $A\sim|Q_{0}|^{2/3}$. Such an exponent evokes some type of hydrodynamic Kolmogorov turbulence, but we lack a theoretical explanation and scaling range is too narrow. Finally, for high-flux runs, we only measure $A$ for the large-$k$ thermodynamic part because the low-$k$ condensate spectrum range is too short for a reliable fitting. Here we see $A=T\sim|Q_{0}|^{1/2}$, which fulfills our prediction for the temperature scaling of the three-wave process.
Figure 3: (a): Percentages of energy components in terms of flux for the incompressible kinetic energy $E_{\rm kinI}$, compressible kinetic energy $E_{\rm kinC}$, quantum energy $E_{\rm q}$, and internal energy $E_{\rm int}$, respectively; (b): “Equation of state” plot covering the weak, intermediate, and strong flux regimes.
@@ Acknowledgements.
This work was funded by the Simons Foundation Collaboration grant Wave Turbulence (Award ID 651471). This work was provided with computing HPC and storage resources of IDRIS and CINES under the allocations A0102A12494 and A0172A14637 made by GENCI on the Jean Zay (SKL partition) and Adastra (GENOA partition) supercomputers. This work was granted access to the OPAL infrastructure from Université Côte d’Azur, supported by the French government, through the UCAJEDI Investments in the Future project managed by the National Research Agency (ANR) under Reference No. ANR-15-IDEX-01.
@@ Appendix C Method to generate the EoS figure
Suppose the spectrum is of the form $n=A\,k^{-x}$, we compute the values of $A$ and $x$ by fitting the power-law range of the numerically observed spectrum with $\ln{n}=\ln{A}-x\ln{k}$. Figure S2 presents how we fit spectra for the strong turbulence cases.
@@
The fitting exponents corresponding to Figure 3 of the main text is shown in Figure S3.
Figure S3: The fitting exponents correspond to Figure 3 of the main text. The theoretical predictions for weak wave turbulence, critical balance, and the Bogoliubov spectrum are 7/3-7/3, 4-4, and 2-2, respectively.